OpenStudy (melstutes):

solve by substitution 3x-2y = 7 x+3y=-5

OpenStudy (anonymous):

3x -2y = 7 ---------(1) x + 3y = -5 x = -5-3y --------(2) Put second in first 3(-5-3y) -2y =7 solve for y then put value of y in 2nd eq

OpenStudy (melstutes):

y=-2?

OpenStudy (anonymous):

yes.